OpenStudy (anonymous):

Apply the translations y = f(x + 5) - 4 to f(x) = (x - 2)^2

OpenStudy (anonymous):

Help a person out..

hartnn (hartnn):

your next maths questions should go in 'Mathematics' section, there you'll get help quickly. to get f(x+5) from f(x), just replace 'x' in f(x) by 'x+5' so, f(x+5) = (x+5-2)^2 = (x+3)^2 then y= f(x+5)-4 = (x+3)^2 -4 if you want you can simplify that to get the form y= ax^2+bx+c